An industry is comprised of five firms. Sales of the four largest firms are $1,000,000, $500,000, $400,000, and $200,000. If the C4 ratio is 96 percent, then what is the HHI?

  • 2,755.

  • 3,046.

  • None of the options.

  • 1,810.

  • 5,017.
Answer #1

C4 = Total sales of four firms / Total sales by all five firms

= (1000000 + 500000 + 400000 + 200000) / (1000000 + 500000 + 400000 + 200000 + x)

= 2100000 / (2100000 + x)

As per given condition

0.96 = 2100000 / (2100000 + x)

(2100000 + x) = 2100000 / 0.96 = 2187500

x = 2187500 - 2100000 = 87500

HHI = 100^2 * [(1000000 / 2187500)^2 + (500000 / 2187500)^2 + (400000 / 2187500)^2 + (200000 / 2187500)^2 + (87500 / 2187500)^2 ]

= 0.3046 * 10000

= 3046

Second option is correct answer